# Disentangled Latent Energy-Based Style Translation: An Image-Level Structural MRI Harmonization Framework

### v1

@@ 5.5 Computational Cost Evaluation
To quantitatively assess the enhancements in computational and time efficiency, we compare our DLEST framework with three GAN-based SOTA methods across several metrics. These include the number of trainable parameters, total floating point operations (FLOPs) per forward pass, training time to convergence on the OpenBHB training dataset, testing time on the SRPBS dataset, and the necessity for retraining when applied to an unseen dataset. As demonstrated in Table 4, our model, encompassing both the SIG and SST modules, contains the fewest parameters, requires significantly fewer FLOPs per forward pass, and generally requires less training time. Furthermore, in comparison to ImUnity, which can also generalize to unseen data without retraining, our DLEST framework requires less time to fine-tune solely the SST module. These results collectively suggest that our DLEST can achieve highly efficient and generalizable MRI harmonization.
@@ 5.6 Limitations and Future Work
While extensive experiments demonstrate that the proposed DLEST is competitive over SOTA methods, it is important to acknowledge some limitations. First, our current implementation focuses on 2D-level MRI slices, potentially overlooking volumetric variations, and can only synthesize discontinued MRI slices instead of the whole volume. One possible solution is to extend our implementation to 2.5D, incorporating a conditioning mechanism in the autoencoder to generate MRI slices by conditioning on adjacent slices (Peng et al., 2022). Second, our current autoencoder employs a one-dimensional latent code to achieve the most efficient encoding and style translation. However, the aggressive 1D encoding discards some inherent spatial structures of the original image, leading to some minor loss of anatomical structures during reconstruction. It is interesting to employ 2D latent encoding (Rombach et al., 2022) with a relatively mild compression rate to better preserve anatomical structures of input MRIs, which will be one of our future works. In addition, our current work only uses latent information from images, without considering biological features, such as gender, age, and relevant clinical diagnostic scores. As our further work, we plan to incorporate biological variables as condition to improve fidelity of generated images.
Table 4: Time and computation cost. For DLEST, $a+b$ denotes the number for SIG $+$ SST. Test time is the cost of generalizing on an unseen dataset. M: Million; GMac: Giga multiply–accumulate operations; H: Hour.
Method
##\##Parameters (M)
FLOPs (GMac)
Training Time (H)
Test Time (H)
Need Re-training
CycleGAN
$28.3$
$1794.1$
$29.1$
$8.0$
$Yes$
StyleGAN
$161.3$
$4865.3$
$53.8$
$9.1$
$Yes$
ImUnity
$671.0$
$246.2$
$25.3$
$1.5$
$No$
DLEST (Ours)
$21.7$+$5.3$
$2.1$+$0.3$
$20.0$+$0.4$
$0.4$
$No$

### v2

@@ 5.5 Computational Cost Evaluation
To quantitatively assess the enhancements in computational and time efficiency, we compare our DLEST framework with three GAN-based SOTA methods across several metrics. These include the number of trainable parameters, total floating point operations (FLOPs) per forward pass, training time to convergence on the OpenBHB training dataset, testing time on the SRPBS dataset, and the necessity for retraining when applied to an unseen dataset. As demonstrated in Table 3, our model, encompassing both the SIG and SST modules, contains the fewest parameters, requires significantly fewer FLOPs per forward pass, and generally requires less training time. Furthermore, in comparison to ImUnity, which can also generalize to unseen data without retraining, our DLEST framework requires less test time to fine-tune solely the SST module. These results collectively suggest that our DLEST can achieve highly efficient and generalizable MRI harmonization.
@@ 5.6 Limitations and Future Work
While extensive experiments demonstrate that the proposed DLEST is competitive over SOTA methods, it is important to acknowledge some limitations. First, our current implementation focuses on 2D-level MRI slices, potentially overlooking volumetric variations, and can only synthesize separate MRI slices instead of the whole volume. One possible solution is to extend our implementation to 2.5D, incorporating a conditioning mechanism in the autoencoder to generate MRI slices by conditioning on adjacent slices [28]. Second, our current autoencoder employs a one-dimensional latent code to achieve the most efficient encoding and style translation. However, the aggressive 1D encoding discards some inherent spatial structures of the original image, leading to some minor loss of anatomical structures during reconstruction. It is interesting to employ 2D latent encoding [31] with a relatively mild compression rate to better preserve anatomical structures of input MRIs, which will be one of our future works. In addition, our current work only uses latent information from images, without considering biological features, such as gender, age, and relevant clinical diagnostic scores. As another further work, we plan to incorporate biological variables as a condition to improve the fidelity of generated images.